# Discrimination performance in illness-death models with interval-censored disease data

### v1

@@ 6 Discussion
This article studied the discrimination performance of a binary time-dependent disease marker in the context of the illness-death model for interval-censored data. A simulation study with several data scenarios was conducted to study four different models: the Cox model with disease as time-dependent marker, the piecewise-constant model implemented in the msm package, the Weibull model and the M-spline model implemented in the SmoothHazard package. Both incident/dynamic and cumulative/dynamic AUC estimates were derived from their transition probabilities and studied. The methods were also applied to a data set of patients surgically treated for soft tissue sarcoma who were scanned for distant metastases at scheduled follow-up visits.
@@
The simulation study showed that the HRs from the piecewise-constant model were less biased than those of the Cox model. The number of patients per data set (1000 vs 2000) did not have a large effect on the estimates of the HR, AUC estimates in incident/dynamic and cumulative/dynamic definition, except for the M-spline model which converged more reliably with large data sets (see Supplementary Material C). The type of censoring scheme did mostly influence AUC estimates based on the Cox model and M-spline model, since hazard estimates are non-parametric and therefore more sensitive at later time points at which fewer events are observed. The Weibull model demonstrated the best performance; however, it had an inherent advantage, as the simulated data followed a Weibull distribution. In practice a Weibull distribution may not be a good fit to the data, as shown in the sof tissue sarcoma application. The M-spline model showed a good performance when estimating the incident/dynamic and cumulative/dynamic AUC however was not always able to converge and provide AUC estimates. AUC estimates based on the piecewise-constant model in the incident/dynamic definition had less bias than those based on the Cox model for scenarios with large spacing between follow-up visits and in the cumulative/dynamic definition they had the largest bias of all methods. The spacing of follow-up visits at which the disease state was observed did have a large effect on estimates of the incident/dynamic AUC, particularly for the Cox model. In practice, depending on how reasonable the assumption of Weibull distributed transition hazards is, one could choose to estimate the AUCs based on the Weibull or M-spline model.
@@
Implicitly, it was assumed that the visiting process did not depend on the state (diseased, non-diseased). In clinical practice however, this may not always be a reasonable assumption. Patients may visit the clinic earlier if they have complaints that are related to being in the diseased state. The effect of an informative visiting process could be a subject of future research.
@@
The performed simulations examined the effect of an interval-censored binary disease marker. Future research could explore the discriminatory performance of an interval-censored, time-dependent covariate with more than two possible values, i.e., a subject that can transition between multiple disease states. This would require estimation methods for general interval-censored multi-state models, which have been developed only in recent years (Machado and van den Hout, , 2018; Gu et al., , 2024; Gomon and Putter, , 2024). Another topic of future research could be to investigate a different definition of the incident/dynamic AUC. At time $t$ it was defined as the probability that the individual that dies at time $t$ was diseased at time $t$; however one could also consider an incident/dynamic AUC in which the disease state is evaluated at an earlier time point $s$, where $s<t$ (Zheng and Heagerty, , 2004).
@@
Acknowledgments. Prof.dr. Michiel van de Sande and the PERsonalized SARcoma Care (PERSARC) study group are gratefully acknowledged for making the soft tissue sarcoma data set available.
@@
Supplementary Material
A Derivation of time-dependent AUC for a binary longitudinal marker
A.1 Incident/dynamic AUC
Let $i,j$ be individuals, $X_{i}(t),X_{j}(t)$ their binary markers at time $t$, and $T_{i}^{D}$ and $T_{j}^{D}$ their death times. The incident/dynamic AUC is defined as
$\text{AUC}^{\text{I/D}}(t)=$ $\Pr\left(X_{i}(t)>X_{j}(t)\mid T_{i}^{D}=t,T_{j}^{D}>t\right)+0.5\Pr\left(X_{i}(t)=X_{j}(t)\mid T_{i}^{D}=t,T_{j}^{D}>t\right)$
$=$ $\Pr\left(X_{j}(t)=0\mid T_{j}^{D}>t\right)\Pr\left(X_{i}(t)=1\mid T_{i}^{D}=t\right)$
$+0.5\cdot\big{[}\Pr\left(X_{j}(t)=0\mid T_{j}^{D}>t\right)\Pr\left(X_{i}(t)=0\mid T_{i}^{D}=t\right)+$
$\hskip 42.67912pt\Pr\left(X_{j}(t)=1\mid T_{j}^{D}>t\right)\Pr\left(X_{i}(t)=1\mid T_{i}^{D}=t\right)\big{]}$
$=$ $\left[1-\Pr\left(X_{j}(t)=1\mid T_{j}^{D}>t\right)\right]\Pr\left(X_{i}(t)=1\mid T_{i}^{D}=t\right)$
$+0.5\cdot\big{\{}\left[1-\Pr\left(X_{j}(t)=1\mid T_{j}^{D}>t\right)\right]\left[1-\Pr\left(X_{i}(t)=1\mid T_{i}^{D}=t\right)\right]+$
$\hskip 45.52458pt\Pr\left(X_{j}(t)=1\mid T_{j}^{D}>t\right)\Pr\left(X_{i}(t)=1\mid T_{i}^{D}=t\right)\big{\}}$
$=$ $\left[1-\pi_{1}(t)\right]p(t)+0.5\left\{\left[1-\pi_{1}(t)\right]\left[1-p(t)\right]+\pi_{1}(t)p(t)\right\}$
$=$ $p(t)-\pi_{1}(t)p(t)+0.5\left[1-p(t)-\pi_{1}(t)+\pi_{1}(t)p(t)+\pi_{1}(t)p(t)\right]$
$=$ $p(t)-\pi_{1}(t)p(t)+0.5-0.5p(t)-0.5\pi_{1}(t)+\pi_{1}(t)p(t)$
$=$ $0.5+0.5\left[p(t)-\pi_{1}(t)\right],$
where
$\pi_{1}(t)=\Pr\left(X_{i}(t)=1\mid T_{i}^{D}>t\right)\quad\text{and}\quad p(t)=\Pr\left(X_{i}(t-)=1\mid T_{i}^{D}=t\right).$
A.2 Cumulative/dynamic AUC
Let $i,j$ be individuals, $X_{i}(s),X_{j}(s)$ their binary markers at time $s$, and $T_{i}^{D}$ and $T_{j}^{D}$ their death times. The cumulative/dynamic AUC is then
$\text{AUC}^{\text{C/D}}(s,t)=$ $\Pr\left(X_{i}(s)>X_{j}(s)\mid s<T_{i}^{D}\leq t,T_{j}^{D}>t\right)+0.5\Pr\left(X_{i}(s)=X_{j}(s)\mid s<T_{i}^{D}\leq t,T_{j}^{D}>t\right)$
$=$ $\Pr\left(X_{j}(s)=0\mid T_{j}^{D}>t\right)\Pr\left(X_{i}(s)=1\mid s<T_{i}^{D}\leq t\right)$
$+0.5\cdot\big{[}\Pr\left(X_{j}(s)=0\mid T_{j}^{D}>t\right)\Pr\left(X_{i}(s)=0\mid s<T_{i}^{D}\leq t\right)+$
$\hskip 42.67912pt\Pr\left(X_{j}(s)=1\mid T_{j}^{D}>t\right)\Pr\left(X_{i}(s)=1\mid s<T_{i}^{D}\leq t\right)\big{]}$
$=$ $\left[1-\Pr\left(X_{j}(s)=1\mid T_{j}^{D}>t\right)\right]\Pr\left(X_{i}(s)=1\mid s<T_{i}^{D}\leq t\right)$
$+0.5\cdot\big{\{}\left[1-\Pr\left(X_{j}(s)=1\mid T_{j}^{D}>t\right)\right]\left[1-\Pr\left(X_{i}(s)=1\mid s<T_{i}^{D}\leq t\right)\right]+$
$\hskip 45.52458pt\Pr\left(X_{j}(s)=1\mid T_{j}^{D}>t\right)\Pr\left(X_{i}(s)=1\mid s<T_{i}^{D}\leq t\right)\big{\}}$
$=$ $\left[1-\pi_{1}(s,t)\right]p(s,t)+0.5\left\{\left[1-\pi_{1}(s,t)\right]\left[1-p(s,t)\right]+\pi_{1}(s,t)p(s,t)\right\}$
$=$ $p(s,t)-\pi_{1}(s,t)p(s,t)+0.5\left[1-p(s,t)-\pi_{1}(s,t)+\pi_{1}(s,t)p(s,t)+\pi_{1}(s,t)p(s,t)\right]$
$=$ $p(s,t)-\pi_{1}(s,t)p(s,t)+0.5-0.5p(s,t)-0.5\pi_{1}(s,t)+\pi_{1}(s,t)p(s,t)$
$=$ $0.5+0.5\left[p(s,t)-\pi_{1}(s,t)\right],$
where
$\pi_{1}(s,t)=\Pr\left(X_{j}(s)=1\mid T_{j}^{D}>t\right)\quad\text{and}\quad p(s,t)=\Pr\left(X_{i}(s)=1\mid s<T_{i}^{D}\leq t\right).$
B Weibull illness-death model
The transition hazards from state $i$ to state $j$ are set to be Weibull distributed and defined as follows:
$\lambda_{hl}(t)=$ $\alpha_{hl}kt^{k-1},$ (19)
where $k$ is the common shape parameter and $\alpha_{hl}$ are transition-specific rate parameters. Let
$S_{0}(t)=$ $\exp\left\{-(\alpha_{01}+\alpha_{02})t^{k}\right\}$
and
$S_{1}(t)=$ $\exp\left\{-\alpha_{12}t^{k}\right\}.$
Then, the transition probabilities are equal to
$P_{00}(u,t)=$ $\frac{S_{0}(t)}{S_{0}(u)},$
$P_{11}(u,t)=$ $\frac{S_{1}(t)}{S_{1}(u)},$
$P_{01}(u,t)=$$\begin{cases}\frac{\alpha_{01}}{\alpha_{01}+\alpha_{02}-\alpha_{12}}\left(\frac{S_{1}(t)}{S_{1}(u)}-\frac{S_{0}(t)}{S_{0}(u)}\right)&\text{if }\alpha_{01}+\alpha_{02}-\alpha_{12}\neq 0\\ \alpha_{01}\left(\frac{S_{1}(t)}{S_{1}(u)}t^{k}-\frac{S_{0}(t)}{S_{0}(u)}u^{k}\right)&\text{otherwise (note: }S_{1}(t)=S_{0}(t)),\end{cases}$
$P_{02}^{0}(u,t)=$ $\frac{\alpha_{02}}{\alpha_{01}+\alpha_{02}}\left(1-\frac{S_{0}(t)}{S_{0}(u)}\right),$
$P_{02}^{1}(u,t)=$$\begin{cases}\frac{\alpha_{01}}{\alpha_{01}+\alpha_{02}}\left(1-\frac{S_{0}(t)}{S_{0}(u)}\right)-\frac{\alpha_{01}}{\alpha_{01}+\alpha_{02}-\alpha_{12}}\left(\frac{S_{1}(t)}{S_{1}(u)}-\frac{S_{0}(t)}{S_{0}(u)}\right)&\text{if }\alpha_{01}+\alpha_{02}-\alpha_{12}\neq 0\\ \frac{\alpha_{01}}{\alpha_{01}+\alpha_{02}}\left(1-\frac{S_{0}(t)}{S_{0}(u)}\right)-\alpha_{01}\frac{S_{0}(t)}{S_{0}(u)}\left(t^{k}-u^{k}\right)&\text{otherwise},\end{cases}$
$P_{02}(u,t)=$ $P_{02}^{0}(u,t)+P_{02}^{1}(u,t)=$
$=$ $1-\frac{\alpha_{02}-\alpha_{12}}{\alpha_{01}+\alpha_{02}-\alpha_{12}}\cdot\frac{S_{0}(t)}{S_{0}(u)}-\frac{\alpha_{01}}{\alpha_{01}+\alpha_{02}-\alpha_{12}}\cdot\frac{S_{1}(t)}{S_{1}(u)},$
and
$P_{12}(u,t)=$ $1-\frac{S_{1}(t)}{S_{1}(u)},$
where $P_{hl}(u,t)$ is the conditional probability of being in state $j$, $(j=0,1,2)$ at time $t$ given in state $i$, $(i=0,1)$ at time $u$.
These transition probabilities can be used to calculate the time-specific incident/dynamic and cumulative/dynamic AUC using Manuscript Equations (9) and (15), respectively.
C Simulation study: Results for all scenarios
Table C.1: Bias, empirical standard error (SE), and root mean square error (RMSE) for the effect of disease (yes vs no) for the Cox and piecewise constant (PW-const) models under scenarios A to L in Manuscript Table 1.
Scenario Model Mean($\beta$) exp(mean($\beta$)) SE($\beta$) Bias($\beta$) RMSE($\beta$)
Truth 2.42 11.20
A Cox ROC 2.35 10.44 0.09 -0.07 0.11
Cox prob 2.35 10.44 0.09 -0.07 0.11
PW-const 2.43 11.32 0.09 0.01 0.09
B Cox ROC 2.29 9.91 0.10 -0.12 0.16
Cox prob 2.29 9.91 0.10 -0.12 0.16
PW-const 2.48 11.90 0.10 0.06 0.12
C Cox ROC 2.24 9.37 0.11 -0.18 0.21
Cox prob 2.24 9.37 0.11 -0.18 0.21
PW-const 2.41 11.11 0.12 -0.01 0.12
D Cox ROC 2.35 10.45 0.09 -0.07 0.11
Cox prob 2.35 10.45 0.09 -0.07 0.11
PW-const 2.45 11.59 0.08 0.03 0.09
E Cox ROC 2.31 10.05 0.09 -0.11 0.14
Cox prob 2.31 10.05 0.09 -0.11 0.14
PW-const 2.50 12.19 0.10 0.08 0.13
F Cox ROC 2.25 9.47 0.10 -0.17 0.20
Cox prob 2.25 9.47 0.10 -0.17 0.20
PW-const 2.43 11.34 0.11 0.01 0.11
G Cox ROC 2.34 10.43 0.07 -0.07 0.10
Cox prob 2.34 10.43 0.07 -0.07 0.10
PW-const 2.42 11.30 0.06 0.01 0.06
H Cox ROC 2.29 9.91 0.07 -0.12 0.14
Cox prob 2.29 9.91 0.07 -0.12 0.14
PW-const 2.47 11.88 0.07 0.06 0.09
I Cox ROC 2.24 9.38 0.08 -0.18 0.20
Cox prob 2.24 9.38 0.08 -0.18 0.20
PW-const 2.41 11.09 0.08 -0.01 0.08
J Cox ROC 2.34 10.40 0.06 -0.07 0.10
Cox prob 2.34 10.40 0.06 -0.07 0.10
PW-const 2.44 11.52 0.06 0.03 0.07
K Cox ROC 2.30 10.00 0.06 -0.11 0.13
Cox prob 2.30 10.00 0.06 -0.11 0.13
PW-const 2.49 12.10 0.07 0.08 0.10
L Cox ROC 2.24 9.40 0.07 -0.17 0.19
Cox prob 2.24 9.40 0.07 -0.17 0.19
PW-const 2.42 11.24 0.08 0.00 0.08
Abbreviations: $\beta$, coefficient of time-dependent variable in Cox model or PW-const model; Cox ROC estimates are obtained via the risksetAUC function from the risksetROC package; Cox prob estimates are based on transition probabilities of Cox model through mstate package.
Table C.2: Number of invalid estimations of $\mbox{AUC}^{I/D}$ and $\mbox{AUC}^{C/D}$ from 1 year based on $n_{sim}=1000$ data sets for the M-spline model under scenarios A to L.
Scenario
A B C D E F G H I J K L
invalid $\text{AUC}^{I/D}$ 379 442 555 17 34 88 201 292 378 1 7 19
invalid $\text{AUC}^{C/D}$ 381 444 560 17 34 89 201 296 381 1 7 19
Table C.3: Bias, empirical standard error (SE), and root mean square error (RMSE) for time-specific incident/dynamic AUC at 1, 3, 5 years from the Cox model under scenarios A to F estimated via the risksetAUC function from the risksetROC R-package.
Scenario Model Bias SE RMSE Bias SE RMSE Bias SE RMSE
A Cox ROC -0.04 0.02 0.04 -0.02 0.01 0.03 -0.02 0.01 0.02
B Cox ROC -0.07 0.02 0.08 -0.04 0.01 0.04 -0.03 0.01 0.04
C Cox ROC -0.21 0.00 0.21 -0.07 0.02 0.08 -0.06 0.02 0.06
D Cox ROC -0.04 0.02 0.04 -0.02 0.01 0.02 -0.02 0.01 0.02
E Cox ROC -0.07 0.02 0.08 -0.04 0.01 0.04 -0.03 0.02 0.03
F Cox ROC -0.21 0.00 0.21 -0.07 0.02 0.08 -0.06 0.02 0.06
Abbreviations: $\text{AUC}^{\text{I/D}}(t)$, incident/dynamic AUC at year $t$
Table C.4: Estimated bias, empirical standard error (SE), and root mean square error (RMSE) for time-specific incident/dynamic AUC at 1, 3, 5 years under scenarios G to L.
Scenario Model Bias SE RMSE Bias SE RMSE Bias SE RMSE
G Cox ROC -0.04 0.01 0.04 -0.02 0.01 0.02 -0.02 0.01 0.02
Cox prob -0.04 0.01 0.04 -0.02 0.01 0.02 -0.02 0.01 0.02
PW-const -0.05 0.01 0.05 -0.03 0.01 0.03 -0.02 0.01 0.03
Weibull 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
M-spline 0.02 0.02 0.03 0.00 0.02 0.02 -0.01 0.02 0.03
H Cox ROC -0.07 0.01 0.08 -0.04 0.01 0.04 -0.03 0.01 0.03
Cox prob -0.07 0.01 0.08 -0.04 0.01 0.04 -0.03 0.01 0.03
PW-const -0.06 0.01 0.06 -0.04 0.01 0.04 -0.03 0.01 0.03
Weibull -0.01 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
M-spline 0.01 0.02 0.02 -0.01 0.02 0.02 -0.01 0.03 0.03
I Cox ROC -0.21 0.00 0.21 -0.07 0.01 0.07 -0.06 0.01 0.06
Cox prob -0.21 0.00 0.21 -0.07 0.01 0.07 -0.06 0.01 0.06
PW-const -0.09 0.01 0.09 -0.06 0.01 0.06 -0.05 0.01 0.05
Weibull -0.01 0.02 0.02 0.00 0.01 0.01 0.00 0.01 0.01
M-spline -0.04 0.03 0.05 0.00 0.02 0.02 0.00 0.03 0.03
J Cox ROC -0.04 0.01 0.04 -0.02 0.01 0.02 -0.02 0.01 0.02
Cox prob -0.04 0.01 0.04 -0.02 0.01 0.02 -0.02 0.01 0.02
PW-const -0.05 0.01 0.05 -0.03 0.01 0.03 -0.02 0.01 0.03
Weibull 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
M-spline 0.02 0.01 0.03 -0.01 0.02 0.02 -0.01 0.02 0.03
K Cox ROC -0.07 0.01 0.07 -0.04 0.01 0.04 -0.03 0.01 0.03
Cox prob -0.07 0.01 0.07 -0.04 0.01 0.04 -0.03 0.01 0.03
PW-const -0.06 0.01 0.06 -0.04 0.01 0.04 -0.03 0.01 0.03
Weibull -0.01 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
M-spline 0.01 0.02 0.02 -0.01 0.02 0.02 0.00 0.02 0.03
L Cox ROC -0.21 0.00 0.21 -0.07 0.01 0.07 -0.06 0.01 0.06
Cox prob -0.21 0.00 0.21 -0.07 0.01 0.07 -0.06 0.01 0.06
PW-const -0.09 0.01 0.09 -0.06 0.01 0.06 -0.05 0.01 0.05
Weibull -0.01 0.02 0.02 0.00 0.01 0.01 0.00 0.01 0.01
M-spline -0.04 0.03 0.05 0.00 0.02 0.02 0.00 0.03 0.03
Abbreviations: $\text{AUC}^{\text{I/D}}(t)$, incident/dynamic AUC at year $t$; PW-const, piecewise-constant model; Cox ROC estimates are obtained via the risksetAUC function from the risksetROC package; Cox prob estimates are based on transition probabilities of Cox model through mstate package.
Figure C.1: Estimated time-specific incident/dynamic AUC for scenario D (3 months; left panels), B (E months; middle panels) and F (12 months; right panels) using different models (Cox, PW-const, Weibull, M-spline). The x-axis represents time $t$ in years; the y-axis represents $\widehat{\text{AUC}}^{\text{I/D}}(t)$. The blue line in each panel represents the true values over time. Estimates for Cox ROC (first row) are based on risksetAUC function for Cox model. Estimates for Cox prob are based on transition probabilities of Cox model via mstate package.
Figure C.2: Estimated time-specific incident/dynamic AUC for scenario G (3 months; left panels), H (E months; middle panels) and I (12 months; right panels) using different models (Cox, PW-const, Weibull, M-spline). The x-axis represents time $t$ in years; the y-axis represents $\widehat{\text{AUC}}^{\text{I/D}}(t)$. The blue line in each panel represents the true values over time. Estimates for Cox ROC (first row) are based on risksetAUC function for Cox model. Estimates for Cox prob are based on transition probabilities of Cox model via mstate package.
Figure C.3: Estimated time-specific incident/dynamic AUC for scenario J (3 months; left panels), K (E months; middle panels) and L (12 months; right panels) using different models (Cox, PW-const, Weibull, M-spline). The x-axis represents time $t$ in years; the y-axis represents $\widehat{\text{AUC}}^{\text{I/D}}(t)$. The blue line in each panel represents the true values over time. Estimates for Cox ROC (first row) are based on risksetAUC function for Cox model. Estimates for Cox prob are based on transition probabilities of Cox model via mstate package.
Table C.5: Estimated bias, empirical standard error (SE), and root mean square error (RMSE) for time-specific cumulative/dynamic AUC for prediction time 1, 3, 5 years and prediction window of 5 years under scenarios G to L.
Scenario Model Bias SE RMSE Bias SE RMSE Bias SE RMSE
G Cox 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
PW-const -0.02 0.00 0.02 -0.01 0.01 0.01 -0.01 0.01 0.01
Weibull 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.01 0.01
M-spline 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
H Cox 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
PW-const -0.04 0.00 0.04 -0.02 0.01 0.02 -0.02 0.01 0.02
Weibull 0.00 0.00 0.01 0.00 0.01 0.01 0.00 0.01 0.01
M-spline 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
I Cox 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
PW-const -0.05 0.00 0.05 -0.04 0.01 0.04 -0.03 0.01 0.03
Weibull -0.01 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
M-spline 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
J Cox 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
PW-const -0.02 0.00 0.02 -0.01 0.01 0.01 -0.01 0.01 0.01
Weibull 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.00
M-spline 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
K Cox 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
PW-const -0.04 0.00 0.04 -0.02 0.01 0.03 -0.02 0.01 0.02
Weibull 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.01 0.01
M-spline 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
L Cox 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
PW-const -0.05 0.00 0.05 -0.04 0.01 0.04 -0.03 0.01 0.03
Weibull -0.01 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
M-spline 0.00 0.01 0.01 0.00 0.01 0.01 0.00 0.01 0.01
Figure C.4: Estimated time-specific cumulative/dynamic AUC for scenario D (3 months; left panels), E (6 months; middle panels) and F (12 months; right panels) using different models (Cox, PW-const, Weibull, M-spline). The x-axis represents the prediction time $t$ in years. The prediction window is set to 5 years, so the y-axis represents $\widehat{\text{AUC}}^{\text{C/D}}(t,t+5)$. The blue line in each panel represents the true values over time.
Figure C.5: Estimated time-specific cumulative/dynamic AUC for scenario G (3 months; left panels), H (6 months; middle panels) and I (12 months; right panels) using different models (Cox, PW-const, Weibull, M-spline). The x-axis represents the prediction time $t$ in years. The prediction window is set to 5 years, so the y-axis represents $\widehat{\text{AUC}}^{\text{C/D}}(t,t+5)$. The blue line in each panel represents the true values over time.
Figure C.6: Estimated time-specific cumulative/dynamic AUC for scenario J (3 months; left panels), K (6 months; middle panels) and L (12 months; right panels) using different models (Cox, PW-const, Weibull, M-spline). The x-axis represents the prediction time $t$ in years. The prediction window is set to 5 years, so the y-axis represents $\widehat{\text{AUC}}^{\text{C/D}}(t,t+5)$. The blue line in each panel represents the true values over time.

### v2

@@ 6 Discussion
This article studied the discrimination performance of a binary time-dependent disease marker in the context of the illness-death model for interval-censored data. A simulation study with several data scenarios was conducted to study four different models: the Cox model with disease as time-dependent marker, the piecewise-constant model implemented in the msm package, the Weibull model and the M-spline model implemented in the SmoothHazard package. Both incident/dynamic and cumulative/dynamic AUC estimates were derived from their transition probabilities and studied. The methods were also applied to a data set of patients surgically treated for soft tissue sarcoma who were scanned for distant metastases at scheduled follow-up visits. In both cases, covariates were not included in the models, but they could be incorporated on a transition-specific basis, allowing each transition $h\rightarrow l$ to have distinct covariate effects, modeled multiplicatively on the hazard as $\exp(\boldsymbol{\beta}^{T}_{hl}\boldsymbol{Z}_{hl})$.
@@
The simulation study showed that the HRs from the piecewise-constant model were less biased than those of the Cox model. The number of patients per data set (400 vs 1000 vs 2000) did not have a large effect on the estimates of the HR, AUC estimates in incident/dynamic and cumulative/dynamic definition, except for the M-spline model which converged more reliably with large data sets (see Supplementary Material C). The type of censoring scheme did mostly influence AUC estimates based on the Cox model and M-spline model, since hazard estimates are non-parametric and therefore more sensitive at later time points at which fewer events are observed. The Weibull model demonstrated the best performance; however, it had an inherent advantage, as the simulated data followed a Weibull distribution. In practice a Weibull distribution may not be a good fit to the data, as shown in the soft tissue sarcoma application. The M-spline model showed a good performance when estimating the incident/dynamic and cumulative/dynamic AUC however was not always able to converge and provide AUC estimates. AUC estimates based on the piecewise-constant model in the incident/dynamic definition had less bias than those based on the Cox model for scenarios with large spacing between follow-up visits and in the cumulative/dynamic definition they had the largest bias of all methods. The spacing of follow-up visits at which the disease state was observed did have a large effect on estimates of the incident/dynamic AUC, particularly for the Cox model. The cumulative/dynamic AUC depends solely on transition probabilities, whereas the incident/dynamic AUC also depends on the underlying transition hazards. Consequently, discrimination measured by the cumulative/dynamic AUC is largely insensitive to differences in baseline hazard specification. This explains why the considered methods do not generally outperform the Cox-based approach for cumulative/dynamic AUC, particularly at follow-up visit times when the Cox model exhibits lower bias.
In practice, depending on how reasonable the assumption of Weibull distributed transition hazards is, one could choose to estimate the AUCs based on the Weibull or M-spline model. However, it is crucial to consider that convergence of M-spline models can be affected by smoothing parameter selection or numerical instability. Automatic smoothing parameter selection via cross-validation, while generally effective, can be computationally intensive, and manual tuning is often challenging in complex datasets and not always feasible in simulation studies. In our application, rescaling the time variable enabled convergence under automatic smoothing parameter selection, highlighting the importance of careful numerical considerations in spline-based modeling. Specifically, extreme values or a wide range of time points can lead to poorly scaled basis functions and penalty terms, complicating optimization and potentially preventing convergence.
The choice between incident/dynamic and cumulative/dynamic discrimination metrics depends on the specific perspective of interest, yet both are scientifically relevant. Incident/dynamic AUC is most appropriate when the focus is on the marker’s ability to identify, at a given time, which individuals who are still at risk are most likely to experience death shortly thereafter. This can be particularly informative for time-specific decision-making, such as screening or monitoring. In contrast, cumulative/dynamic AUC evaluates the marker’s ability to discriminate which individuals who are still at risk at a given time are most likely to experience death before a given time horizon, thereby capturing its long-term prognostic capacity. Since clinical and research objectives may require either short-term or cumulative assessments, considering both approaches provides a more comprehensive evaluation of marker performance.
@@
A limitation of multi-state models, such as the illness-death model, is that small sample sizes (e.g., 100 or 200 subjects) often lead to unreliable estimates, even when event times are exactly observed, due to the complexity of modelling multiple transitions and the potential sparsity of events. This issue is further amplified in the presence of interval censoring. Consequently, caution is warranted when applying the proposed methods in such settings. Moreover, it was implicitly assumed that the visiting process did not depend on the state (diseased, non-diseased). In clinical practice however, this may not always be a reasonable assumption. Patients may visit the clinic earlier if they have complaints that are related to being in the diseased state. The effect of an informative visiting process could be a subject of future research.
@@
The performed simulations examined the effect of an interval-censored binary disease marker. Future research could explore the discriminatory performance of an interval-censored, time-dependent covariate with more than two possible values, i.e., a subject that can transition between multiple disease states. This would require estimation methods for general interval-censored multi-state models, which have been developed only in recent years (Machado and van den Hout, , 2018; Gu et al., , 2024; Gomon and Putter, , 2025). Another topic of future research could be to investigate a different definition of the incident/dynamic AUC. At time $t$ it was defined as the probability that the individual that dies at time $t$ was diseased at time $t$; however one could also consider an incident/dynamic AUC in which the disease state is evaluated at an earlier time point $s$, where $s<t$ (Zheng and Heagerty, , 2004).
@@
Supplementary Material
Supplementary Material is available online in Statistical Methods in Medical Research along with the published article (DOI: 10.1177/09622802251412855).
Funding. A.J. R.-B. was supported by the KWF Kankerbestrijding grant UL2015-8028. M.S. was supported by the KWF Kankerbestrijding grant 2023-3 DEV/15461. The funding sources had no involvement in study design, in the collection, analysis and interpretation of the data, in writing of the report, and in the decision to submit the article for publication.
Acknowledgments. The simulation study was performed using the compute resources from the Academic Leiden Interdisciplinary Cluster Environment (ALICE) provided by Leiden University. Prof.dr. Michiel van de Sande and the PERsonalized SARcoma Care (PERSARC) study group are gratefully acknowledged for making the soft tissue sarcoma data set available.
@@
Data availability statement. The soft-tissue sarcoma data are not publicly available due to privacy restrictions. The R code used to generate the simulated data sets as described in Section 4.1 and perform the simulation study is provided at {http://github.com/mspreafico/auc-idmIC}.